# Exploring the Design Space of Diffusion Bridge Models via Stochasticity Control

### v1

@@ Abstract
Diffusion bridge models effectively facilitate image-to-image (I2I) translation by connecting two distributions. However, existing methods overlook the impact of noise in sampling SDEs, transition kernel, and the base distribution on sampling efficiency, image quality and diversity. To address this gap, we propose the Stochasticity-controlled Diffusion Bridge (SDB), a novel theoretical framework that extends the design space of diffusion bridges, and provides strategies to mitigate singularities during both training and sampling. By controlling stochasticity in the sampling SDEs, our sampler achieves speeds up to 5×5$\times$5 × faster than the baseline, while also producing lower FID scores. After training, SDB sets new benchmarks in image quality and sampling efficiency via managing stochasticity within the transition kernel. Furthermore, introducing stochasticity into the base distribution significantly improves image diversity, as quantified by a newly introduced metric.
@@ 1 Introduction
Denoising Diffusion Models (DDMs) create a stochastic process to transition Gaussian noise into a target distribution (Song & Ermon, 2019; Ho et al., 2020; Song et al., 2020). Building upon this, diffusion bridge-based models (DBMs) have been developed to transport between two arbitrary distributions, $\pi_{T}$ and $\pi_{0}$, including Bridge Matching (Peluchetti, 2023), Flow Matching (Lipman et al., 2022), and Stochastic Interpolants (Albergo et al., 2023). Compared to DDMs, DBMs offer greater versatility for tasks such as I2I translation (Linqi Zhou et al., 2023; Liu et al., 2023). This advantage arises because using a Gaussian prior often fails to incorporate sufficient knowledge about the target distribution.
In general, there are two primary design philosophies for DBMs. The first involves deriving a pinned process (Yifeng Shi et al., 2023) from a given reference process (e.g., Brownian motion) via Doob’s $h$-transform, and then constructing a bridge to approach it  (Linqi Zhou et al., 2023; Peluchetti, 2023). The second regime aims to directly design a bridge based on a specified transition kernel (Lipman et al., 2022; Albergo et al., 2023). While the former also results in a transition kernel, the mean and variance in the kernel are coupled, which limits the design flexibility for possible bridges. In this work, we follow the second fashion and further propose the Stochasticity Control (SC) mechanism, which facilitates easier tuning and leads to enhanced performance across a variety of tasks. Our main contributions are as follows:
  • We introduce the Stochasticity-controlled Diffusion Bridge (SDB), a generalized framework that adopts a transition kernel-based design philosophy to elucidate the design space of DBMs. Notably, this framework not only encompasses other mainstream DBMs such as DDBM (Linqi Zhou et al., 2023) and I2SB (Liu et al., 2023), but also DDMs like EDM (Karras et al., 2022), as detailed in Table 1.
  • A Stochasticity Control (SC) mechanism is proposed by adding noise into the base distribution, designing a noise schedule for the transition kernel, and regulating the drift term in the sampling SDEs. In addition, we explore score reparameterization and the discretization schemes of sampling SDEs to mitigate singularity during training and sampling. These combined strategies lead to significant improvements in training stability, sampling efficiency, output quality, and conditional diversity.
  • Experimental results show that our sampler operates 5×5$\times$5 × faster than the DDBM sampler and achieves a lower FID score using the same pretrained models. When trained from scratch, our model sets a new benchmark for image quality, requiring only $5$ function evaluations to reach an FID of $0.89$. Furthermore, by introducing noise into the base distribution, we significantly enhance the diversity of synthetic images, resulting in a greater variety of colors and textures.
Notations Let $\pi_{T}$, $\pi_{0}$, and $\pi_{0T}$ represent the base distribution, the target distribution, and the joint distribution of them respectively. $\pi_{\mathrm{cond}}$ and $\pi_{\mathrm{data}}$ represent the distributions of the input and output data. Let $p$ be the distribution of a diffusion process; we denote its marginal distribution at time $t$ by $p_{t}$, the conditional distribution at time $t$ given the state at time $s$ by $p_{t|s}$, and the distribution at time $t$ given the states at times 00 and $T$ by $p_{t|0T}$, i.e., the transition kernel of a bridge.
@@ 2 Background
2.1 Denoising diffusion models
Denoising diffusion models map target distribution $\pi_{0}$ into a base distribution $\pi_{T}$ by define a forward process on the time-interval $[0,T]$:
$d\mathbf{X}_{t}=\bar{f}_{t}\mathbf{X}_{t}dt+\bar{g}_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{0}\sim\pi_{0},$ (1)
where $\bar{f}_{t},\bar{g}_{t}:[0,T]\to\mathbb{R}$ is the scalar-valued drift and diffusion term, $\mathbf{X}_{0}\in\mathbb{R}^{d}$ is drawn from the target distribution $\pi_{0}$, $\mathbf{W}_{t}$ is a $d$-dimensional Wiener process. To sample from the target distribution $\pi_{0}$, the generative model is given by the reverse SDE or ODE (Song et al., 2020):
$d\mathbf{X}_{t}=\left[\bar{f}_{t}\mathbf{X}_{t}-\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log q_{t}(\mathbf{X}_{t})\right]dt+\bar{g}_{t}d\mathbf{W}_{t},\quad\mathbf{X}_{T}\sim\pi_{T},$ (2)
$d\mathbf{X}_{t}=\left[\bar{f}_{t}\mathbf{X}_{t}-\frac{1}{2}\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log q_{t}(\mathbf{X}_{t})\right]dt,\quad\mathbf{X}_{T}\sim\pi_{T},$ (3)
where $q_{t}$ denotes the marginal distribution of this process. The score function $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ is approximated using a neural network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, which can be learned by the score-matching loss:
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{x}_{t}\sim p_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0}),\mathbf{x}_{0}\sim\pi_{0},t\sim\mathcal{U}{(0,T)}}\left[\omega(t)\left\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})\right\|^{2}\right],$ (4)
where $q_{t|0}$ is the analytic forward transition kernel and $\omega(t)$ is a positive weighting function.
@@ 2.2 Denoising Diffusion Bridge Models
DDBMs (Linqi Zhou et al., 2023) extend diffusion models to translate between two arbitrary distributions $\pi_{0}$ and $\pi_{T}$ given samples from them. Consider a reference process in Eq. (1) with transition kernel $q_{t|0}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};a_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I})$, this process can be pinned down at an initial and terminal point $\mathbf{x}_{0},\mathbf{x}_{T}$. Under mild assumptions, the pinned process is given by Doob’s $h$-transform (Rogers & Williams, 2000):
$\mathrm{d}\mathbf{X}_{t}=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}\mathbf{W}_{t},\quad\mathbf{X}_{0}=\mathbf{x}_{0},$ (5)
where $\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}\mid\mathbf{X}_{t})=\frac{(a_{t}/a_{T})\mathbf{x}_{T}-\mathbf{X}_{t}}{\sigma_{t}^{2}(\mathrm{SNR}_{t}/\mathrm{SNR}_{T}-1)}$ and $\mathrm{SNR}:=a_{t}^{2}/\sigma_{t}^{2}$ (Linqi Zhou et al., 2023). The marginal density of process (5) serves as transition kernel and is given by $p(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0}{+}\beta_{t}\mathbf{x}_{T},\gamma_{t}^{2}\mathbf{I})$, where $\alpha_{t}=a_{t}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}}),\beta_{t}=\frac{a_{t}}{a_{T}}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}},\gamma_{t}^{2}=\sigma_{t}^{2}(1{-}\frac{\mathrm{SNR}_{T}}{\mathrm{SNR}_{t}})$.
To sample from the conditional distribution $p(\mathbf{x}_{0}|\mathbf{x}_{T})$, we can solve the reverse SDE or probability flow ODE from $t=T$ to $t=0$:
$\mathrm{d}\mathbf{X}_{t}$ =$=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}(\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})-\nabla_{\mathbf{X}_{t}}\log p_{t|T}(\mathbf{X}_{t}|\mathbf{x}_{T}))\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}\mathbf{W}_{t},\mathbf{X}_{T}=\mathbf{x}_{T}$ (6)
$\mathrm{d}\mathbf{X}_{t}$ =$=\{\bar{f}_{t}\mathbf{X}_{t}+\bar{g}_{t}^{2}(\nabla_{\mathbf{X}_{t}}\log p_{T|t}(\mathbf{x}_{T}|\mathbf{X}_{t})-\frac{1}{2}\nabla_{\mathbf{X}_{t}}\log p_{t|T}(\mathbf{X}_{t}|\mathbf{x}_{T}))\}\mathrm{d}t,\quad\mathbf{X}_{T}=\mathbf{x}_{T}.$ (7)

### v2

@@ Abstract
Diffusion bridge models and stochastic interpolants enable high-quality image-to-image (I2I) translation by creating paths between distributions in pixel space. However, the proliferation of techniques based on incompatible mathematical assumptions have impeded progress. In this work, we unify and expand the space of bridge models by extending Stochastic Interpolants (SIs) with preconditioning, endpoint conditioning, and an optimized sampling algorithm. These enhancements expand the design space of diffusion bridge models, leading to state-of-the-art performance in both image quality and sampling efficiency across diverse I2I tasks. Furthermore, we identify and address a previously overlooked issue of low sample diversity under fixed conditions. We introduce a quantitative analysis for output diversity and demonstrate how we can modify the base distribution for further improvements.
@@ 1 Introduction
Denoising Diffusion Models (DDMs) and flow matching create a stochastic process to transition Gaussian noise into a target distribution (song2019generative, ; ho2020denoising, ; song2020score, ; lipman2022flow, ). Building upon this, diffusion bridge-based models (DBMs) have been developed to transport between two arbitrary distributions, $\pi_{T}$ and $\pi_{0}$, including I2SB (liu2023i2sb, ), DSBM (yifengshiDiffusionSchrOdinger2023, ), DDBM (linqizhouDenoisingDiffusionBridge2023, ), DBIM zheng2024diffusion , Bridge Matching (peluchetti2023non, ). DBMs achieve superior image quality in I2I translation compared to DDMs (linqizhouDenoisingDiffusionBridge2023, ; liu2023i2sb, ; albergo2023stochasticdatacoupling, ), primarily because the distance between source and target image distributions is typically smaller than that between Gaussian and target distributions.
While DBMs like DDBM (yifengshiDiffusionSchrOdinger2023, ), DBIM (zheng2024diffusion, ), and I2SB (liu2023i2sb, ) achieve state-of-the-art FID scores in image-to-image translation, they suffer from limited design flexibility, constrained bridge path formulations, and complex parameter tuning. In contrast, Stochastic Interpolants (SIs) (albergo2023stochastic, ; albergo2023stochasticdatacoupling, ) offer a simpler and more flexible framework, but they have yet to integrate practical advances from recent diffusion bridge models, such as preconditioning. Besides, SIs require training two separate models, unlike the more efficient single-model setup in DDBM. Table 1 summarizes the key characteristics of these methods, highlighting that their complementary strengths had not yet been unified.
Another overlooked issue stemming from restrictive design choices in previous bridge models is the lack of diversity in outputs. While some image translation tasks are one-to-one, we find that in one-to-many translation tasks, like black and white edges to color images, previous methods produce limited variation in colors and textures. We refer to this as the conditional diversity problem and show that our approach leads to significant improvements.
Figure 1: The design space of bridge paths and samplers.
DDBM DBIM DSBM SI ECSI (ours)
Endpoint conditioning
Uncoupled parameters
Extensive bridge paths
Extensive samplers
Preconditioning
Modified base density
Table 1: Characteristics of different bridge models.
Figure 2: Samples for I2I translation with our ECSI models: Deblurring, Depth-RGB, and Edges to Handbags. For each pair of images, we show the input image (upper) and the output image (bottom).
Our main contributions are as follows:
  • We propose Endpoint-Conditioned Stochastic Interpolants (ECSI), which extend stochastic interpolants by incorporating endpoint conditioning and preconditioning. Previous bridge methods artificially coupled unrelated aspects of the transition kernel. ECSI introduces a decoupled parametrization that expands and simplifies the design space for bridge paths and samplers. To further improve sampling quality and efficiency, we develop a novel noise control scheme and an efficient sampling algorithm.
  • We identify a previously overlooked issue: the low diversity of outputs conditioned on fixed source images. To address this, we propose modifying the base distribution. Furthermore, to quantitatively evaluate conditional output diversity, we introduce a new metric—Average Feature Diversity (AFD).
  • Experimental results demonstrate our model’s state-of-the-art performance in both image quality and sampling speed across various I2I tasks, including deblurring, edges-to-handbags translation, and depth-to-RGB conversion. Notably, for handbag generation, our approach yields significantly more diverse outputs with varied colors and textures.
@@ 2 Background
Notations Let $\pi_{T}$, $\pi_{0}$, and $\pi_{0T}$ represent the base distribution, the target distribution, and the joint distribution of them respectively. $\pi_{\mathrm{cond}}$ and $\pi_{\mathrm{data}}$ represent the distributions of the input and output data. Let $p$ be the distribution of a diffusion process; we denote its marginal distribution at time $t$ by $p_{t}$, the conditional distribution at time $t$ given the state at time $s$ by $p_{t|s}$, and the distribution at time $t$ given the states at times 00 and $T$ by $p_{t|0,T}$, i.e., the transition kernel of a bridge.
@@ 2.1 Denoising Diffusion Bridge Models
DDBMs (linqizhouDenoisingDiffusionBridge2023, ) extend diffusion models to translate between two arbitrary distributions $\pi_{0}$ and $\pi_{T}$ given samples from them. Consider a reference process given by:
$d{X}_{t}=\bar{f}_{t}{X}_{t}dt+\bar{g}_{t}d{W}_{t},$ (1)
whose transition kernel is given by $q_{t|0}({x}_{t}|{x}_{0})=\mathcal{N}({x}_{t};a_{t}{x}_{0},\sigma_{t}^{2}\mathbb{I})$. This process can be conditioned (or "pinned") at both an initial point ${x}_{0}$ and a terminal point ${x}_{T}$ to construct a diffusion bridge. Under mild assumptions, the pinned process is given by Doob’s $h$-transform (rogers2000diffusions, ):
$\mathrm{d}{X}_{t}=\{\bar{f}_{t}{X}_{t}+\bar{g}_{t}^{2}\nabla_{{X}_{t}}\log p_{T|t}({x}_{T}|{X}_{t})\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}{W}_{t}$ (2)
where $\nabla_{{X}_{t}}\log p_{T|t}({x}_{T}\mid{X}_{t})=\frac{(a_{t}/a_{T}){x}_{T}-{X}_{t}}{\sigma_{t}^{2}(\mathrm{SNR}_{t}/\mathrm{SNR}_{T}-1)}$ and $\mathrm{SNR}_{t}:=a_{t}^{2}/\sigma_{t}^{2}$ (linqizhouDenoisingDiffusionBridge2023, ). Eq. 2 is a stochastic process that transport from $p_{0}=\pi_{0}$ and $p_{t}=\pi_{t}$, which is a valid bridge process. To sample from the conditional distribution $p({x}_{0}|{x}_{T})$, we can solve the reverse SDE or probability flow ODE from $t=T$ to $t=0$:
$\mathrm{d}{X}_{t}$ =$=\{\bar{f}_{t}{X}_{t}+\bar{g}_{t}^{2}(s-h)\}\mathrm{d}t+\bar{g}_{t}\mathrm{d}{W}_{t},$ (3)
$\mathrm{d}{X}_{t}$ =$=\{\bar{f}_{t}{X}_{t}+\bar{g}_{t}^{2}(s-\frac{1}{2}h)\}\mathrm{d}t,$ (4)